OpenStudy (anonymous):

FOR THIS I GOT D. Am i right? Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = -1 x = 0 x = 2 There is no x value as there is no solution. x can be any value as there is an infinite number of solutions.

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

2x + y = 1 4x + 2y = -1 These are two parallel lines, they never meet.

OpenStudy (anonymous):

So i am right?

ganeshie8 (ganeshie8):

you can figure that they're parallel by dividing the second equation by 2

ganeshie8 (ganeshie8):

dividing the second equation by 2, and subtracting gives : 2x + y = 1 2x + y = -1/2 ----------------- 0 = 3/2 which is FALSE.

ganeshie8 (ganeshie8):

so, there will not be any solutions.

ganeshie8 (ganeshie8):

and nope, D is not correct as it says there are infinitely many solutions.

OpenStudy (anonymous):

So C

ganeshie8 (ganeshie8):

yup ! C is correct

OpenStudy (anonymous):

Can you help me with2 more that i haven't figured out?

ganeshie8 (ganeshie8):

shoot...

OpenStudy (anonymous):

Solve the following system of equations: x - 2y = 14 x + 3y = 9 (1, 12) (-1, -12) (12, -1) (12, 1)

OpenStudy (anonymous):

Solve the system of equations. x + 3y = -1 2x + 2y = 6 (-4, 1) (2, -1) (4, -1) (5, -2)